It was clear last night, after two weeks of cloud, and not as cold as the last clear evening. I opened the observatory at about 8:00 with a temp of -5, planning to acquire one or more colours of RGB for the B33 project I started a few weeks ago.
Not a very successful session – I ended up gathering a small amount of Red data, and several items for my to-do list.
I had to re-do the sky pointing model since I had to manualy slew the mount back to rest position last time. Something strange with the sky model, but I didn’t take the time to track it down – the first couple of reference go-tos were not far off, but the next few were considerably worse. Odd, perhaps the time or location have become messed up or, more likely, some kind of drag or slippage in the clutches. To be followed up when it’s warmer.
I framed the same B33 image as last time, and set up guiding. Next problem: guiding calibration kept failing, reporting “no motion in Y-axis”. Manually sending guiding pulses to the scope, I eventually tracked this down to a fault in the autoguider cable. Wiggling the cable and re-seating the connectors, I eventually got it working.
To-do list: make spare guider cables, and make them longer so there is room to secure them against the connection surface.
Notes for the operations manual: when guiding calibration fails, check the control cable as a first step.
With guiding finally working, I started a 30-minute acquisition for Red data. Focused with the Red filter in place and scheduled 6 5-minute subs. I went indoors, monitoring the control computer via VNC, to stay warm while this was happening.
Guiding wasn’t very good. Not sure what has changed since last excellent session, but I was getting quite good X-axis guiding (small-fractional-pixel errors on the graph) but Y-axis guiding was swinging around quite a lot, with a distinct sine wave of nearly a pixel-width of correction. I’m wondering if PEC (which was on) is somehow out of sync, or if mount parameters have changed substantially because of the change in temperature.
To-do list: do another PEC training run.
To-do list: in warmer weather, re-check the worm gear clearance, backlash, and stiction.
With the Y-axis guiding aggressiveness turned up a bit, I got half-pixel guiding that looked like it would be adequate, and let the exposures run. 5 of the 6 worked ok. The 6th failed, saying “guide star faded”. Going back outside, I see that what actually happened was that tracking stopped because the scope hit the RA tracking limit, which seems to be set too high.
To-do list: reset the RA tracking limits to proper values.
I packed everything up, dropping the laptop on the deck and breaking a corner of the plastic in the process, and closed down for the night.
To-do list: install an inexpensive permanent computer in the observatory to avoid having to transport laptop in and out each night.
Having a quick look at the acquired data, the dust spot blooming is *really* bad – more than I should be expecting simple flat frames to correct. Because it was also this bad, and almost identical, on Luminance test frames, I assume the problem is on the chip or its protective window, not on the filters.
To-do list: bring camera and filters indoors for a careful cleaning.
Finally, I tried having a quick look at the acquired data. To my surprise, MaximDL refuses to use the dark frames and flat frames I gathered last week to calibrate the new images. Although I know that wouldn’t have been ideal calibration (temperature having changed) I still expected it to work, and was careful not to move the camera so the flats would be useable. No idea why it won’t accept calibration this way, and the error message is not very informative.
To-do list: research the error message and figure out why the calibration frames weren’t acceptable.
So, next evening, with clean chip and filters, I’ll try this all again. Good thing I tell myself that my hobby is the process, not the results.